### v1

@@ I Introduction
The conjecture’s appeal lies not only in its theoretical depth but also in the computational challenges it presents. These challenges have spurred the development of algorithms to analyze the stopping time of numbers in the Collatz sequence—the number of steps required for a given number to reach 1 [3]. Efficient computation of stopping times is crucial for verifying the conjecture across large numeric ranges and for exploring its potential implications in algorithmic design.
This paper introduces a novel algorithm for calculating the stopping time of numbers in the Collatz sequence. The proposed algorithm achieves significant improvements in computational efficiency by leveraging structural patterns within the Collatz tree to minimize redundant operations. Experimental evaluations demonstrate that the algorithm consistently outperforms existing state-of-the-art methods, achieving computational efficiency improvements of approximately 28% across various input sizes, ranging from small to extremely large numbers.
@@ II Literature Review
The literature on the Collatz conjecture spans diverse perspectives and approaches, broadly categorized into three main areas: (A) theoretical insights and structural analysis, (B) computational methods and algorithmic efficiency, and (C) practical applications and implications. Theoretical studies focus on structural behaviors, probabilistic models, and generalized formulations of the conjecture. Computational advancements explore algorithmic designs that push the limits of numeric verification. Finally, practical applications leverage the conjecture’s iterative properties for real-world problems, such as cryptography, blockchain consensus, and digital watermarking. The following sections provide a detailed review of these categories.
@@ A. Theoretical Insights and Structural Analysis
An approximation method analyzes stopping times in the Collatz sequence by defining transformations as $T(m)=m/2$ for even $m$ and $T(m)=(3m+1)/2$ for odd $m$. This approach establishes probabilistic bounds, showing that average stopping times for most integers fall within a logarithmic framework. The model provides lower bounds, contributing insights into the asymptotic behavior of stopping times if the conjecture holds [5].
@@
An extensive review of computational and theoretical efforts to find Collatz counterexamples shows no success in uncovering a counterexample. The analysis suggests that iterative search reduction might contribute to proof efforts or reveal that the conjecture could be an unprovable truth in mathematics, deepening the study of this open problem [8].
A statistical analysis assesses the probability of infinite stopping times in the Collatz sequence, which would contradict the conjecture. The results show that this likelihood is astronomically low, likened to selecting a specific atom in the universe, thus supporting the assumption that all sequences eventually reach 1 [9].
@@ B. Computational Methods and Algorithmic Efficiency
In addition to bitwise techniques, a set of novel algorithms and theorems analyze the structural relationships within Collatz sequences. These tools include a directed graph visualization, an iteration analyzer, and a peak value finder that collectively explore the relationship between natural numbers, peak values, and iteration counts. By revealing these structural elements, these computational tools provide insights that could contribute to proving the conjecture [11].
Automata-based verification offers further computational efficiency by processing transformations in batches rather than individually. This method utilizes a “Collatz tree” organized by residue classes to enable direct processing of certain values without full transformations. The automata design outputs segments of Collatz transformations in one step, effectively grouping repetitive patterns and significantly optimizing the verification process [12].
@@ III Methodology
This section details the methodology for developing and analyzing an efficient algorithm to compute stopping times in the Collatz sequence. It begins with an exploration of structural patterns in the Collatz tree, which serve as the foundation for algorithmic optimization. The proposed algorithm is then described, followed by an analysis of its computational efficiency and correctness. A comparison with state-of-the-art methods provides context for the improvements achieved by the new approach. Verification of the algorithm ensures its reliability in adhering to the properties of the Collatz sequence.
@@ III-A The Collatz Tree and Observed Patterns
The Collatz tree is a hierarchical representation of the reverse transformations of the Collatz sequence, starting from $1$. As illustrated in Fig. 1, the structural patterns of the tree provide a foundation for developing efficient algorithms by identifying key paths and simplifying redundant computations.
@@ III-B Algorithm Development for Calculating Stopping Time
Handling Even Numbers: If the current number is even, it belongs to a branch rooted at an odd number. To find the root of the current branch, determine the largest power of 2 that divides the current number with a reminder of zero. This can be computed efficiently using the bitwise operation:
@@
$largest\_power=20,480\land-20,480=4,096$
$exponent=\log_{2}(4,096)=12$
$stopping\_time=0+12=12$
$n=20,480/4,096=5$
@@
$largest\_power=16\land-16=16$
$exponent=\log_{2}(16)=4$
$stopping\_time=13+4=17$
$n=16/16=1$
@@
$stopping\_time=17+1=18$
Hence, the stopping time for $n=20,480$ is $18$. As we will see in the following sections, even the current state-of-the-art algorithms require 18 iterations to compute this result. However, the proposed algorithm optimizes the process by leveraging structural patterns in the Collatz tree, significantly reducing computational redundancy.
@@ III-C Analysis of Proposed Algorithm
  • One iteration moves 3 to the parent sub-branch starting with $5$. ($3*3+1=10$)
@@
  • One iteration moves 5 to the base branch. ($5*3+1=16$)
@@
  • $k$ iterations are required to reduce the numbers in each sub-branch to their root odd numbers
  • $k$ iterations are required to traverse up the tree to the parent sub-branches
  • One iteration is required at the base branch to reduce the number to $1$
@@
  • The formula encapsulates the traversal of both even and odd numbers in the Collatz sequence.
@@
Confirmation with $12k+$ 1: In the best case ($k=0$), the formula:
Total $\text{Total Iterations}=2k+1=1,$
shows the number is reduced to $1$ in $O(1)$ time without traversing sub-branches.
Time Complexity: Each operation ($\log_{2}(n)$, bitwise AND, division) is $O(1)$, making the overall best-case complexity $\mathbf{O(1)}$.
@@
Visualization: Figure 2 provides a visual representation of the iterations across different input sizes. The scatter plots demonstrate how iterations vary with input size, revealing the logarithmic growth of iterations with respect to the input numbers. Each subplot corresponds to a selected input range, offering a detailed view of the scaling behavior.
@@
Average-Case Time Complexity: To derive the average-case complexity, we analyzed the number of iterations for exponentially increasing input sizes. The observed average number of iterations for input sizes $10,000$, $20,000$, …, $5,120,000$ indicated slow, logarithmic growth.
@@
Assuming the average number of iterations follow the form:
@@
The formula for the average number of iterations as a function of $n$ is therefore:
$f(n)=4.65\cdot\log_{2}(n)-4.91.$
Order of Growth The derived formula confirms that the average number of iterations grows logarithmically with the input size. Therefore, the average-case complexity is $\mathbf{O(\log(n))}$, where n is the input size. This result aligns with the dominance of even reductions and the relatively shallow depth of sub-branches observed in the Collatz tree structure.
@@
$f(n)=28\cdot\log_{2}(n)-180.17.$
Order of Growth: The logarithmic relationship confirms that the worst-case complexity is $\mathbf{O(\log(n))}$, where n is the input size.
@@ III-D Verification of The Proposed Algorithm
The experiment utilized OpenCL to leverage parallel computing capabilities, thereby significantly reducing computation time. Both the brute-force and proposed algorithms were executed in parallel for each number in the range $1$ to $1,000,000,000$. For each input, the stopping times computed by both algorithms were compared to ensure consistency and correctness across the entire range.
@@ V-B Significance
VI Conclusion
This research introduces a novel algorithm for calculating the stopping times in the Collatz sequence, focusing on computational efficiency and scalability. By leveraging structural patterns in the Collatz tree, the proposed algorithm minimizes redundant operations and achieves consistent improvements of approximately 28% over state-of-the-art methods. Unlike prior approaches, the algorithm efficiently handles extremely large numbers without requiring advanced techniques such as memoization, showcasing its robustness and practicality.
The findings demonstrate the algorithm’s potential for extending the verification of the Collatz conjecture to larger numeric ranges. While this research prioritizes algorithmic design and optimization, large-scale verification of the conjecture remains an important direction for future work. Additionally, the algorithm’s low space complexity and scalability open avenues for applications in other areas of computational mathematics, such as sequence analysis, cryptography, and random number generation.
Future research will explore further optimizations, including the integration of memoization and parallel computing, to enhance performance and enable the verification of significantly larger ranges. These advancements will contribute to the ongoing computational exploration of the Collatz conjecture and similar problems in number theory.

### v2

@@ I Introduction
The enduring fascination with the Collatz conjecture stems from its ability to bridge elementary number theory and advanced computational techniques, making it both accessible and challenging. The conjecture’s appeal lies not only in its theoretical depth but also in the computational challenges it presents. These challenges have spurred the development of algorithms to analyze the stopping time of numbers in the Collatz sequence—the number of steps required for a given number to reach 1 [3]. A precise understanding of stopping times is critical for large-scale verification efforts and has implications for optimizing computational strategies.
This paper introduces a novel algorithm for calculating the stopping time of numbers in the Collatz sequence. The proposed algorithm achieves significant improvements in computational efficiency by leveraging structural patterns within the Collatz tree to minimize redundant operations. Key optimizations include identifying and exploiting repetitive patterns within the tree, allowing the algorithm to bypass unnecessary computations. Experimental evaluations demonstrate that the algorithm consistently outperforms existing state-of-the-art methods, achieving computational efficiency improvements of approximately 28% across various input sizes, ranging from small to extremely large numbers.
@@ II Literature Review
The literature on the Collatz conjecture spans diverse perspectives and approaches, broadly categorized into three main areas: (A) theoretical insights and structural analysis, (B) computational methods and algorithmic efficiency, and (C) practical applications and implications. Theoretical studies examine the conjecture’s mathematical foundations and the structural intricacies of its behavior. These studies aim to uncover patterns, propose generalizations, and assess probabilistic models. Computational advancements explore algorithmic designs that push the limits of numeric verification. Finally, practical applications leverage the conjecture’s iterative properties for real-world problems, such as cryptography, blockchain consensus, and digital watermarking. The following sections provide a detailed review of these categories.
@@ A. Theoretical Insights and Structural Analysis
An approximation method analyzes stopping times in the Collatz sequence by defining transformations as $T(m)=m/2$ for even $m$ and $T(m)=(3m+1)/2$ for odd $m$. This method introduces probabilistic bounds, showing that average stopping times fall within a logarithmic framework and providing insights into potential asymptotic behaviors of the conjecture. The model provides lower bounds, contributing insights into the asymptotic behavior of stopping times if the conjecture holds [5].
@@
An extensive review of computational and theoretical efforts to find Collatz counterexamples shows no success in uncovering a counterexample. The analysis also suggests that iterative reduction strategies might either contribute to proving the conjecture or demonstrate its potential as an unprovable truth in mathematics, enriching its theoretical study [8].
A statistical analysis assesses the probability of infinite stopping times in the Collatz sequence, which would contradict the conjecture. The results indicate that the likelihood of divergence is astronomically low, analogous to selecting a specific atom from the entire observable universe, reinforcing the assumption that all sequences eventually converge to 1 [9].
@@ B. Computational Methods and Algorithmic Efficiency
In addition to bitwise techniques, a set of novel algorithms and theorems analyze the structural relationships within Collatz sequences. These tools include a directed graph visualization, an iteration analyzer, and a peak value finder that collectively explore the relationship between natural numbers, peak values, and iteration counts. Such tools provide computational insights into the conjecture’s complex structure, which could support eventual proof efforts [11].
Automata-based verification offers further computational efficiency by processing transformations in batches rather than individually. This method utilizes a “Collatz tree” organized by residue classes to enable direct processing of certain values without full transformations. By segmenting repetitive patterns, this design significantly optimizes the verification process and provides a scalable framework for future work [12].
@@ III Methodology
This section outlines the methodology for developing and analyzing an efficient algorithm to compute stopping times in the Collatz sequence. It begins with an exploration of structural patterns in the Collatz tree, which serve as the foundation for algorithmic optimization. The proposed algorithm is then described, followed by an analysis of its computational efficiency and correctness. The algorithm’s efficiency is compared with state-of-the-art methods, followed by its verification to ensure adherence to the Collatz sequence properties.
The experiments were conducted on two different hardware configurations. All experiments in subsections III-C, III-D, IV-A, and V were conducted on a MacBook Air equipped with an Apple M2 chip, 8GB of unified memory, and running macOS Sequoia Version 15.2. These experiments utilized Python 3.12.4 and Julia 1.11.2. The experiment in subsection IV-B was conducted on Google Colab using an NVIDIA T4 GPU with Python 3.11.11.
@@ III-A The Collatz Tree and Observed Patterns
The Collatz tree is a hierarchical representation of the inverse transformations of the Collatz sequence, starting from 1. As shown in Fig. 1, the structural patterns of the tree provide a foundation for developing efficient algorithms by identifying key paths and simplifying redundant computations.
@@ III-B Algorithm Development for Calculating Stopping Time
Handling Even Numbers: If the current number is even, it belongs to a branch rooted at an odd number. To find the root of the current branch, determine the largest power of 2 that divides the current number with a remainder of zero. This can be computed efficiently using the bitwise operation:
@@
[…]
$\textnormal{exponent}=\log_{2}(4{,}096)=12$
[…]
$n=\frac{20{,}480}{4{,}096}=5$
@@
[…]
$\textnormal{exponent}=\log_{2}(16)=4$
[…]
$n=\frac{16}{16}=1$
@@
[…]
Hence, the stopping time for $n=20{,}480$ is $18$. While the state-of-the-art algorithms also compute this result in $18$ iterations, the proposed algorithm achieves the same result using only $3$ iterations by directly leveraging structural patterns in the Collatz tree. This optimization significantly reduces computational redundancy and highlights the algorithm’s efficiency.
@@ III-C Analysis of Proposed Algorithm
  • One iteration moves 3 to the parent sub-branch starting with $5$ ($3\times 3+1=10$)
@@
  • One iteration moves 5 to the base branch ($5\times 3+1=16$)
@@
  • $k$ iterations are required to reduce the numbers in each sub-branch to their root odd numbers.
  • $k$ iterations are required to traverse up the tree to the parent sub-branches.
  • One iteration is required at the base branch to reduce the number to $1$.
@@
  • Equation 4 encapsulates the traversal of both even and odd numbers in the Collatz sequence.
@@
Confirmation with $12k+$ 1: In the best case ($k=0$), the total iterations is $12k+1=2(0)+1=$ 1, and this confirms that any number on the base branch reduces to $1$ in $O(1)$, without requiring any traversal through sub-branches.
Time Complexity: Each operation ($\log_{2}(n)$, bitwise AND, division) is performed in $O(1)$, leading to an overall best-case complexity of $\mathbf{O(1)}$.
@@
Visualization: Figure 2, generated with Matplotlib, provides a visual representation of the iterations across different input sizes. The scatter plots demonstrate how iterations vary with input size, revealing the logarithmic growth of iterations with respect to the input numbers. Each subplot corresponds to a selected input range, offering a detailed view of the scaling behavior.
@@
Average-Case Time Complexity: To derive the average-case complexity, we analyzed the number of iterations for exponentially increasing input sizes. The observed average number of iterations for input sizes $10{,}000$ to $5{,}120{,}000$ indicated slow, logarithmic growth.
@@
Assuming the average number of iterations follows the form:
@@
Therefore, the average number of iterations as a function of $n$ is given by:
$f(n)=4.65\cdot\log_{2}(n)-4.91$ (5)
Order of Growth Equation 5 confirms that the average number of iterations grows logarithmically with the input size. Consequently, the algorithm’s average-case complexity is $\mathbf{O(\log(n))}$, where $n$ is the input size. This result aligns with the dominance of even reductions and the relatively shallow depth of sub-branches observed in the Collatz tree structure.
@@
$f(n)=28\cdot\log_{2}(n)-180.17$ (6)
Order of Growth: Equation 6 confirms that the worst-case complexity scales logarithmically with the input size. Therefore, the algorithm’s worst-case complexity is $\mathbf{O(\log(n))}$, where $n$ is the input size.
@@ III-D Verification of The Proposed Algorithm
The experiment utilized OpenCL 1.2 to leverage parallel computing capabilities, thereby significantly reducing computation time. Parallelization was employed solely for accelerating the verification process and was not used in other experiments where the performance of the proposed algorithm was measured. Both the brute-force and proposed algorithms were executed in parallel for each number in the range $1$ to $1,000,000,000$. For each input, the stopping times computed by both algorithms were compared to ensure consistency and correctness across the entire range.
@@ V-B Significance
VI Limitations, Scope, and Future Work
The performance of the proposed algorithm is inherently constrained by the structure of the Collatz tree. As established in Equation 4, the total number of iterations depends on the number of branches, which are determined by the frequency of odd numbers in the sequence. Since all branches start with odd numbers, an increase in their occurrence leads to a higher branch count $k$, increasing the number of iterations. However, experimental results confirm a logarithmic relationship between input size and the number of iterations, even in the worst-case scenario, ensuring that the algorithm maintains scalability and efficiency across a wide range of inputs.
The proposed algorithm is specifically designed for efficient stopping time computations based on the structural properties of the Collatz tree. Given that stopping time calculations are crucial for large-scale verification of the conjecture, this work primarily benefits applications requiring such computations.
While the proposed algorithm efficiently computes stopping times, integrating memoization and parallelization can further enhance performance in large-scale verification processes. Parallelization allows the algorithm to run across multiple CPU or GPU cores, enabling concurrent computation of stopping times, while memoization reduces redundant calculations by storing previously computed values. These optimizations primarily benefit scenarios where multiple numbers need to be processed in bulk.
Future research will focus on large-scale verification of the Collatz conjecture by employing memoization and parallelization to the proposed algorithm. By leveraging these techniques, significantly larger numeric ranges can be verified, further contributing to the computational exploration of the conjecture.
VII Conclusion
This research introduces an optimized algorithm for computing stopping times in the Collatz sequence, leveraging structural properties of the Collatz tree to minimize redundant operations. Experimental evaluations confirm that the proposed algorithm achieves a consistent 28% reduction in computational iterations compared to state-of-the-art methods, demonstrating its efficiency across varying numeric ranges. Unlike prior approaches, the algorithm efficiently handles extremely large numbers without requiring auxiliary optimization techniques such as memoization or parallelization, reinforcing its scalability and robustness.
A comprehensive experimental analysis was conducted to evaluate the algorithm’s efficiency under different input conditions. The results highlight its ability to process small and large random numbers, powers of two, multiples of three, and prime numbers with improved execution times. Furthermore, by analyzing the worst-case and average-case complexity, this study establishes that the total number of iterations follows a logarithmic relationship with input size, ensuring stable performance even for large-scale computations.
Beyond algorithmic optimization, this work provides insights into the structural behavior of the Collatz tree, contributing to a deeper understanding of sequence evolution. While the primary focus remains on optimizing stopping time calculations, the approach has broader implications in computational mathematics, including applications in sequence analysis, cryptography, and large-scale numerical verification.
In summary, this study advances the state of computational approaches to the Collatz conjecture, offering a highly efficient algorithm for stopping time calculations. The proposed optimizations provide a foundation for extending computational verification efforts in number theory while opening new possibilities for algorithmic applications in theoretical and applied mathematics.